OpenStudy (anonymous):

square root of 14s+49 =s-7

OpenStudy (anonymous):

14s + 49 = s^2 - 14s + 49 28s = s^2 s = 28

OpenStudy (anonymous):

square root of t-8 =2t+1

OpenStudy (ybarrap):

sqrt(t-8)=2t+1 t-8 = (2t + 1)^2 t-8 = (4t^2 + 4t + 1) 4t^2 + 3t + 9 = 0 4[(t^2 + 3t/4) + 9/4] = 0 4[(t + 3/8)^2 - 9/64 + 9/4] = 0 4[(t + 3/8)^2 + (- 9 + 9*16)/64] = 0 4[(t + 3/8)^2 + 135/64] = 0 (t + 3/8)^2 = - 135/64 t+3/8 = +/- sqrt(-135/64) t = sqrt(-134)/8 - 3/8 and t = -sqrt(-135)/8 - 3/8 Ans. t = sqrt(-134)/8 - 3/8 and t = -sqrt(-135)/8 - 3/8

OpenStudy (ybarrap):

Note that solutions are not real (i.e. they are complex).